OpenStudy (anonymous):

Find 4 consecutive odd integers where the product of the two smaller numbers is 64 less than the product of the two larger numbers. @johnweldon1993

OpenStudy (johnweldon1993):

Can you apply what I did in your last question to this one?

OpenStudy (anonymous):

idk but it wont be the same answer thought

OpenStudy (anonymous):

cuz they want 4 consecutive for this problem

OpenStudy (johnweldon1993):

that's correct....this time it is odd numbers....we can think of this as, whatever even number you have...when you add an odd number to it...you'll have an odd number right? So here you'll have 2n + 1, 2n + 3, 2n + 5, and 2n + 7 this will give you 4 numbers...now can you use those old steps to solve it?

OpenStudy (johnweldon1993):

Here I'll do the equation again for you and see if you can continue it product of the 2 smaller numbers is 64 less than the product of the 2 bigger numbers (2n + 1)(2n + 3) = (2n + 5)(2n + 7)

OpenStudy (johnweldon1993):

Woops....use this new equation (2n + 1)(2n + 3) = (2n + 5)(2n + 7) - 64

OpenStudy (anonymous):

3,5,7,9 i this the correct answer

OpenStudy (johnweldon1993):

We can check your answer the product of the 2 smaller numbers you have (15) is 64 less than the product of the 2 larger numbers (63) ...so no that cannot be correct...try out my equation

OpenStudy (johnweldon1993):

(2n + 1)(2n + 3) = (2n + 5)(2n + 7) - 64 \[4n^2 + 8n + 3 = 4n^2 + 24n + 35 - 64\] can you continue on?

OpenStudy (anonymous):

can u gib=ve me the answer plz

OpenStudy (johnweldon1993):

I cannot give you the answer...I can show you how to do it again for future reference

OpenStudy (johnweldon1993):

\[8n + 3 = 24n - 29\] \[-16n = -32\] \[n = 2\] now that you know 'n' = 2 you can go back to those original numbers we had 2n + 1, 2n + 3, 2n + 5, 2n + 7 and replace all the 'n' with 2 2(2) + 1 = 5 2(2) + 3 = 7 2(2) + 5 = 9 2(2) + 7 = 11

OpenStudy (johnweldon1993):

now we can check product of 2 smaller (35) is 64 less than product of 2 larger (99) 99 - 35 = 64 so those numbers are correct